OpenStudy (anonymous):

3 log base 5 y - 1og base 25 y = 10 This is a difficult question for me to solve and i hope you would help me thanks.

OpenStudy (anonymous):

I know that log base 25 is actually 1/2 log base 5 y

hartnn (hartnn):

sure, first you need to know this property of log, \[\log_ba= \dfrac{\log a}{\log b}\]

OpenStudy (anonymous):

understood the property

hartnn (hartnn):

so, \(\log_5 y= ...?? \\ \log_{25}y=... ?\)

OpenStudy (anonymous):

log base 5 y = log y / log 5 and log base 25 y = log y/ log 25

hartnn (hartnn):

good :) now \(\log x^n=n \log x \\ so, \log 25=.... ?\)

OpenStudy (anonymous):

2 log 5 sir

hartnn (hartnn):

so, now you can put log y =x and the equation will become ...?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

can you give me the full step solution lol

OpenStudy (anonymous):

i am getting confuse here

hartnn (hartnn):

\(3x/log 5 -x/(2log 5) = 10 \\ \) got this step first ?

hartnn (hartnn):

where x=log y

OpenStudy (anonymous):

keep it going

hartnn (hartnn):

multiply both sides by 2log 5 6x - x = 20 log 5 5x = 20 log 5 got this much ?

OpenStudy (anonymous):

The first step that you posted is that the first step of showing the workings?

hartnn (hartnn):

if you understand how to solve, you will be able to show the workings on your own....

OpenStudy (anonymous):

20 log 5 means 20 log base 5?

hartnn (hartnn):

i have made the base unspecified (when its unspecified, by default its 10), so 20 log 5 means 20 times log of 5 (with the base 10)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i think that is where the problem lies in. if not, i could easily understand your workings.

hartnn (hartnn):

\(\log_ba= \dfrac{\log_{10} a}{\log_{10} b} \\\log _{25}y =\dfrac{\log_{10} y}{\log_{10} 25} \) now ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

\(\log_ba= \dfrac{\log_{10} a}{\log_{10} b} \\\log _{25}y =\dfrac{\log_{10} y}{\log_{10} 25}=\dfrac{\log_{10} y}{2\log_{10} 5}\\\log_{5}y=\dfrac{\log_{10} y}{\log_{10} 5} \\ let \log y=x \\ 3x/\log 5 -x/(2\log 5)=10 \implies 6x-x=20\log_{10}5\\x=4\log_{10}5 \\ \log_{10} y = \log _{10}5^4 \implies y= 5^4\)

OpenStudy (anonymous):

okay so there is only 1 value for x?

hartnn (hartnn):

yes, and so 1 value of y.

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

i want to see whether i got the right answer.